OpenStudy (anonymous):

write an equation for each line. slop= -1; through (-3,5)

Vocaloid (vocaloid):

m = slope (x1,y1) = your point equation: y - y1 = m(x-x1)

OpenStudy (anonymous):

I don't understand?

Vocaloid (vocaloid):

read your problem again

Vocaloid (vocaloid):

m = slope = ?

OpenStudy (anonymous):

so m=-1?

Vocaloid (vocaloid):

yes

Vocaloid (vocaloid):

(x1,y1) = the point given in the problem x1 = ? y1 = ?

OpenStudy (anonymous):

x=-1 y=(-3,5 ?

Vocaloid (vocaloid):



Vocaloid (vocaloid):

x1 = ? y1 = ?

OpenStudy (anonymous):

x1 = -3 y1 = 5

Vocaloid (vocaloid):

good. now we have m = -1 x1 = -3 y1 = 5 now plug in y1, x1, and m into the equation y - y1 = m(x-x1) and tell me what you get

OpenStudy (anonymous):

y-5=-1(X-3) ??

OpenStudy (anonymous):

im sorry, Im not good at math at all.

Vocaloid (vocaloid):

almost x - x1 = x - (-3) = ?

OpenStudy (anonymous):

i still don't know lol